Huron Company produces a commercial cleaning compound known as Zoom. The direct materials and direct labor standards for one unit of Zoom are given below: Standard Quantity or Hours Standard Price or Rate Standard Cost Direct materials 7.30 pounds $ 2.80 per pound $ 20.44 Direct labor 0.60 hours $ 9.00 per hour $ 5.40 During the most recent month, the following activity was recorded: 16,600.00 pounds of material were purchased at a cost of $2.50 per pound. All of the material purchased was used to produce 2,000 units of Zoom. 1,000 hours of direct labor time were recorded at a total labor cost of $12,000. Required: 1. Compute the materials price and quantity variances for the month. 2. Compute the labor rate and efficiency variances for the month. (For all requirements, Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values. Round your intermediate calculations to the nearest whole dollar.
Answer #1

1. Material price variance = (standard price - actual price)*actual quantity

= (2.80-2.50)*16600 = $4980 favorable

Material quantity variance = (2000*7.30 - 16600)*2.80 = $5600 unfavorable

2. Labor rate variance = (9-12000/1000)*1000 = $3000 unfavorable

Labor efficiency variance = (2000*0.60 -1000)*9 = $1800 favorable